OpenStudy (anonymous):

If I=indefinite integral of (sqrt(a^2x^2-b^2)/x) dx and x=c*sec(u) so that the substitution reduces I to b*integral of (tan(u))^2*du. Hence evaluate I.

OpenStudy (anonymous):

HELP?

OpenStudy (anonymous):

Does anyone have a clue?

zepdrix (zepdrix):

\[\large \int\limits \frac{\sqrt{a^2x^2-b^2}}{x}dx\]It's a little bit hard to read in text, does this look correct?

zepdrix (zepdrix):

If you're using Internet Explorer, you'll have trouble viewing the code. Lemme know if it's just a bunch of ugly jumbled up text.

OpenStudy (anonymous):

yes thats correct

zepdrix (zepdrix):

They're just telling you to make a `Trig Substitution` in order to solve this integral. Have you had much work with trig subs yet? This one is a tiny bit tricky.

OpenStudy (anonymous):

we just started trig subs. Im fine with it but I am not successful with this one at all

OpenStudy (anonymous):